THE RAMAN FREQUENCY SHIFTS RELATED TO THE THERMODYNAMIC QUANTITIES CLOSE TO THE PHASE I - PHASE II TRANSITION IN NH_{4}Br

\underline{H. YURTSEVEN}, Department of Physics, Middle East Technical University, Ankara, 06531 - Turkey.

We establish here the modified Pippard relations by relating the specific heat C_{p} to the frequency shifts \frac{1}{ν} (\frac{∂ν}{∂T} )_{P} and also the thermal expansivity α_{p} to the \frac{1}{ν} (\frac{∂ν}{∂P} )_{T} for the ν (140 cm^{−1}) Raman mode of NH_{4}Br close to its phase I- phase II transition. From our linear plots of C_{p} vs. \frac{1}{ν} (\frac{∂ν}{∂T} )_{P} and α_{p} vs. α_{p} to the \frac{1}{ν} (\frac{∂ν}{∂P} )_{T}, we extract the values of the slope DP/dT, which are close to the experimental value of 28.2 bar/^{∘}C near the phase I-phase II transition in NH_{4}Br.